### v3

# RangeSAM: On the Potential of Visual Foundation Models for Range-View represented LiDAR segmentation

Abstract
LiDAR point cloud segmentation is central to autonomous driving and 3D scene understanding. While voxel- and point-based methods dominate recent research due to their compatibility with deep architectures and ability to capture fine-n geometry, they often incur high computational cost, irregular memory access, and limited runtime efficiency due to scaling issues. In contrast, range-view methods, though relatively underexplored - can leverage mature 2D semantic segmentation techniques for fast and accurate predictions. Motivated by the rapid progress in Visual Foundation Models (VFMs) for captioning, zero-shot recognition, and multimodal tasks, we investigate whether SAM2, the current state-of-the-art VFM for segmentation tasks, can serve as a strong backbone for LiDAR point cloud segmentation in the range view representations. We present RangeSAM, to our knowledge, the first range-view framework that adapts SAM2 to 3D segmentation, coupling efficient 2D feature extraction with projection/back-projection to operate on point clouds. To optimize SAM2 for range-view representations, we implement several architectural modifications to the encoder: (1) a novel Stem module that emphasizes horizontal spatial dependencies inherent in LiDAR range images, (2) a customized configuration of Hiera Blocks tailored to the geometric properties of spherical projections, and (3) an adapted Window Attention mechanism in the encoder backbone specifically designed to capture the unique spatial patterns and discontinuities present in range-view pseudo-images. Our approach achieves competitive performance on SemanticKITTI while benefiting from the speed, scalability, and deployment simplicity of 2D-centric pipelines. This work highlights the viability of VFMs as general-purpose backbones for point cloud segmentation and opens a path toward unified, foundation-model-driven LiDAR segmentation. Results lets us conclude that range-view segmentation methods using VFMs leads to promising results.
1 Introduction
Figure 1: Our approach leverages the Segment Anything foundation model (SAM2) [64] on range‐view projections of 3D point clouds: (1) render the point cloud as a range image; (2) apply SAM2 to obtain high‐quality 2D masks; (3) back‐project these masks onto the original point cloud, yielding improved segmentation accuracy.
Reconstructing urban scenes along with its 2D and 3D semantics is an important task in computer vision with numerous applications. In the context of creating a photorealistic simulator for autonomous systems like self-driving vehicles, drones, mobile robots, etc.—it becomes crucial to holistically represent all aspects of the scene relevant to driving. This includes tasks such as synthesizing images from interpolated and extrapolated viewpoints in real-time, reconstructing 2D and 3D semantics, generating depth information, and tracking dynamic objects. Among these perception tasks, semantic segmentation of LiDAR point clouds is fundamental, enabling real-time differentiation between vehicles, road signs, pedestrians, and infrastructure. In recent years, research has focused on voxel- and point-based segmentation methods [57, 94, 95, 113, 34, 78]. While these approaches achieve strong performance, they impose substantial computational and memory costs on large-scale outdoor LiDAR data [63, 93, 103, 88, 83, 60, 23, 80, 61] and are challenged by the irregular, unordered nature of point clouds, which complicates efficient processing and network design. In contrast, range-view segmentation [41, 15, 3, 54, 62], projecting the 3D point cloud into a dense 2D image-like representation offers an attractive alternative. It enables reuse of mature 2D semantic segmentation models, reduces memory footprint, and supports faster inference. Despite these advantages, range-view methods have often been overlooked or deemed inferior due to earlier limitations in handling occlusions, resolution loss, and limited context modeling. Recent advances in deep learning, including attention mechanisms [82], multi-scale fusion strategies [111], and context-aware architectures [53], invite a reassessment of the range-view paradigm.
We propose RangeSAM, which utilizes vision foundation models, specifically Segment Anything Model 2 (SAM2) [64], to segment a range view representation (see Figure 1). By employing the SAM2 encoder, we effectively segment the range view representation, resulting in a refined segmented point cloud. RangeSAM not only inherits the robust zero-shot capabilities of SAM2 for 2D scene understanding but also extends these benefits to the 3D representation through segmentation in the range view representation. Our methodology encompasses a range-projecting pre-processing step that transforms unordered LiDAR scans into a range view representation, followed by an novel multi-component model architecture that incorporates Receptive Field Blocks [46] to improve segmentation accuracy. We also implement effective postprocessing techniques, including label propagation via $k$-NN interpolation, to ensure high-quality segmentation results on dense datasets. The model’s training is guided by a composite loss function that addresses class imbalance and promotes accurate boundary delineation. Our main contributions can be summarized as follows:
  • We introduce RangeSAM, the first LiDAR point cloud segmentation framework that lifts knowledge of SAM2 to segment a point cloud utilizing range-view representation derived from unordered LiDAR scans (see Figure 1).
  • Our multi-component encoder architecture leverages the pretrained Hiera backbone, incorporating a Stem module for tensor transformation, a novel embedding matrix to enhance feature sensitivity, and Hiera blocks that utilize both localized and global Multi-Head Attention mechanisms (see Figure 2).
  • We present the experimental results obtained using our proposed RangeSAM methodology, which shows promising potential on the SemanticKITTI benchmark [5], achieving competitive results that, while not state-of-the-art, validate the viability of our approach. Additionally, we conducted an ablation study to explore how various training and data augmentation strategies affect model performance.
Figure 2: Overview of our SAM2-based point-cloud segmentation model. The stem reshapes range-view point clouds into a tensor suitable for the encoder. The encoder is built from stacked Hiera blocks—each containing a multi-head self-attention module and a feed-forward network [64]. The decoder comprises Receptive Field Blocks [46] (RFB) with LayerNorm [4] and GELU [33], concatenates multi-scale features, and projects them to $N_{\text{classes}}$ while also adding auxiliary head (Aux) on corresponding output.
2 Related Work
Visual Foundation Models
Recent advances in large-scale vision models have revolutionized image downstreaming tasks [106, 97, 47, 7]. Notably, the Segment Anything Model (SAM) series achieved promptable image segmentation with impressive zero-shot transferability [38, 71, 67, 44]. Its follow-up SAM2 extends this capability to videos and more complex scenarios [64, 14, 98, 36, 11, 48, 105]. Further the SAM framework was used to leverage its capabilities from 2D into 3D space [101, 104, 8]. Other vision foundation models include self-supervised transformers like the DINO series [10, 56, 72] and multimodal models like [59, 116] and the EVA series [25, 75, 76, 26], which learn rich visual features at scale. These models have been shown to provide strong generalization and label-efficient learning. For instance, DINOv2 features can be extracted and injected into a LiDAR segmentation network to dramatically improve 3D semantic segmentation on both indoor and outdoor benchmarks [107, 101]. Likewise, CLIPs joint image-text embeddings have been transferred to 3D point cloud understanding to enable zero-shot or few-shot 3D segmentation [13, 59, 108, 77].
2.1 Methods on 3D representations
Point-Serialization Methods: Recent point-based serialization-based approaches [85, 16, 49] diverge from conventional paradigms by imposing structural organization through systematic spatial sorting, transforming unstructured point data into ordered sequences while preserving local coherence. Methods based on PointNet [61, 63, 62, 114] directly process unstructured point sets with permutation-invariant features, though their irregular format poses scalability challenges for large-scale LiDAR data. Transformer-based architectures have addressed this limitation. The Point Transformer (PT) series [57, 94, 95] introduced self-attention with relative positional encoding in PTv1 , PTv2 enhanced efficiency through grouped vector attention and partition-based pooling, while PTv3 integrated efficient attention with self-supervised pretraining [96], achieving state-of-the-art performance on outdoor and indoor benchmark datasets. FlatFormer [49] addresses computational efficiency by flattening point clouds through window-based sorting into equal-sized groups rather than equal-shaped windows, applying grouped self-attention with alternating sorting axes and window shifting. OctFormer [86] employs octree-based structured attention for linear scaling through hierarchical spatial partitioning.
Voxel-Based Methods: Unlike 2D approaches, voxel-based methods [79, 52, 31, 29] work directly in 3D space by partitioning the point cloud data into discrete voxels using cartesian coordinates, followed by the application of 3D convolutional operations for processing. Early works like VoxNet [51] demonstrated dense 3D convolution feasibility but suffered from prohibitive memory consumption. OctNet [68] addressed this through octree-based hierarchical voxelization. The paradigm shift toward sparse representations culminated with MinkowskiNet [18], which became the standard through efficient sparse tensor operations and hash-based coordinate mapping. Subsequent developments [113, 34, 78] further optimized efficiency through cylindrical coordinates and neural architecture search. Additionally, numerous approaches for indoor scene point cloud segmentation [63, 93, 103, 88, 83, 60, 23, 80, 61] utilize architectures inspired by PointNet [61], which operate directly on unstructured 3D point data without requiring preprocessing or spatial transformations, these methods do not readily transfer to outdoor scene point cloud segmentation tasks due to computational challenges arising from the high point density. Recent transformer-based voxel architectures [50, 32, 45] synergistically combine sparse convolution efficiency [19] with self-attentions representational power, capturing long-range dependencies while maintaining computational tractability.
Figure 3: Building blocks of the stem module in RangeSAM
2.2 Methods on Projection-Based representations
Convolutional Neural Networks (CNNs): Projection-based methods transform irregular 3D point clouds into structured 2D representations, like range-, birds-eye- or perspective-view enabling direct application of mature 2D CNN architectures [99, 115, 109, 6, 5, 12, 81, 80, 65, 54, 42, 39, 35, 21]. Early work SqueezeSeg model series [90, 91, 99, 84] pioneered real-time segmentation with lightweight CNNs and Conditional Random Fields (CRF) [22] post-processing. Several semantic segmentation methods, including DarkNetSeg [5], SalsaNext [21], KPRNet [39], RangeNet++ [54], Lite-HDSeg [65], and SqueezeSegV3 [99], employ a two-stage approach: they first convert point cloud data into range images through range projection, then apply U-Net-style architectures [69] to process these projected images. While works based on RangeSeg [15, 87] incorporated multi-scale context modeling for spherical projections.
Vision Transformers (ViTs): ViT architectures have catalyzed significant advances: [3, 102] introduced hybrid CNN-Transformer backbones combining convolutional inductive biases with global attention capabilities benefiting from transfer-learning from 2D datasets like [20], while RangeFormer [41] established a fully transformer-based architecture achieving state-of-the-art performance on several outdoor benchmark datasets [5, 6, 9, 74, 58].
3 Methodology
We introduce our projection-based LiDAR segmentation model leveraging VFMs - RangeSAM. We detail the range-projecting preprocessing, the modifications we contribute to the models achitecture and postprocessing procedures. The overall processing pipeline is shown in Figure 1, while Figure 23 and 4 provides a detailed schematic of the stem module, the SAM2-based encoder, and the decoder components.
3.1 Range Projection
Each LiDAR scan is initially represented as an unordered point set $\mathcal{P}={(x,y,z,f)}$, where $(x,y,z)$ denotes the Cartesian coordinates and $f$ represents auxiliary sensor measurements such as intensity or remission values. Subsequently, the point cloud undergoes transformation to a range view representation via projection onto the sensors spherical coordinate system [54]:
$\theta=\arctan\left(\frac{y}{x}\right),\phi=\arcsin\left(\frac{z}{r}\right),r=\sqrt{x^{2}+y^{2}+z^{2}}$ (1)
Specifically, the 3D points are discretized through rasterization into a 2D cylindrical projection with dimensions $H\times W$:
$\binom{u}{v}=\binom{\frac{1}{2}\left[1-\arctan(y,x)\pi^{-1}\right]w}{\left[1-\left(\arcsin\left(zr^{-1}\right)+\mathfrak{f}_{\mathrm{up}}\right)\mathfrak{f}^{-1}\right]h}$, (2)
where $(u,v)$ represent the projected coordinates of point $p_{n}$ in the range image coordinate system, and $h$ and $w$ denote the image height and width parameters, respectively. $r=\sqrt{x^{2}+y^{2}+z^{2}}$ denotes the range between a point and its sensor, the sensors vertical field of view $\mathfrak{f}$ comprises the sum of ($\mathfrak{f}_{\mathrm{up}}$ ++ $\mathfrak{f}_{\mathrm{down}}$) viewing angles. For our method, we configured the imaging system to operate at the commonly used spatial resolution of $64\times 2048$ pixels [3, 15, 41]. Multiple points projecting to the same pixel are resolved by retaining the minimum-range feature. Unprojected pixels are zero-filled.
Figure 4: Building blocks of the encoder module in RangeSAM
3.2 Model Architecture
SAM2 represents a state-of-the-art VFM trained on the SA-V dataset, constituting the largest video segmentation dataset available at the time of its release [64]. SAM2-UNet [98] has achieved superior performance across multiple segmentation tasks, including camouflaged object detection, salient object detection, and mirror detection, while maintaining architectural simplicity. RangeSAM, adopts a similar architectural paradigm (see Figure 2), incorporating Receptive Field Blocks (RFB)[46] for feature decoding to enhance segmentation performance.
First, we have the stem module (see Figure 3) that maps the input tensors so that it can be utilized by the SAM2 encoder.
Stem: This module progressively transforms input tensors from shape $(B,6,H,W)$ to $(B,96,H,W)$ through a sequential application of linear transformation, layer normalization [4], and GELU activation [33] (see Figure 3). Subsequently, the feature maps are partitioned into overlapping $7\times 7$ patches with unit stride. We omit original SAM2 positional embedding scheme and introduce a novel embedding matrix of dimensions $(4,128)$ to enhance the models sensitivity to horizontal spatial features (see Figure 2).
Encoder: SAM2 utilizes the pretrained Hiera [70, 64] backbone. For point cloud segmentation tasks, we design a multi-component encoder architecture (see Figure 4) comprising the following elements:
Hiera block: Each stage of the backbone consists of a customized number of Hiera blocks, and each block includes two modules: 1) Multi-Head Attention module, in early stages, attention is restricted within a window, while later stages may also use global attention. The operation can be written as
$X_{out}=X+\text{DropPath}(\text{MHA}(\text{LayerNorm}(X)))$ (3)
$\text{MHA}(Q,K,V)=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h})W^{O}$ (4)
where
$\begin{split}\operatorname{head}_{i}=\operatorname{Attention}\left(QW_{Q}^{i},KW_{K}^{i},VW_{V}^{i}\right),\\ \operatorname{Attention}(Q,K,V)=\sigma\left(\frac{QK^{\top}+M}{\sqrt{d_{\text{head }}}}\right)V\end{split}$ (5)
where DropPath represents the stochastic depth regularization technique [43], and LayerNorm denotes the layer normalization method [4], which provides batch size-independent normalization capabilities. 2) Feed-forward network:
$O=\operatorname{DropPath}(\operatorname{MLP}(\operatorname{LayerNorm}(O)))\oplus X_{\text{out }},$ (6)
$F=\operatorname{Lin}\left(\operatorname{GELU}\left(\operatorname{DWConv}2\mathrm{D}_{3\times 3}(\operatorname{GELU}(\mathrm{O}))\right)\right)+\mathrm{O}$ (7)
where $\mathrm{DWConv2D}_{3\times 3}$ denotes the 3×3 depthwise convolution operation that introduces spatial locality inductive bias while maintaining computational efficiency through minimal parameter overhead. The SAM2-tiny backbone architecture comprises four Hierarchical stages containing $[1,2,7,2]$ Hiera blocks respectively. Global attention mechanisms are implemented at blocks $[5,7,9,11]$ to capture long-range spatial dependencies.
Attention Window: In the early stages, Hiera employs local windowed attention with a masking strategy where $M_{ij}=0$ for tokens $i$ and $j$ within the same mask unit, and $M_{ij}=$-$\infty$ otherwise. Later stages utilize global attention with $M=0$. Given the range-view image resolution of 64 × 2048, we propose an asymmetric attention window design that emphasizes horizontal spatial relationships: $8\times 64$ for the first and fourth stages, and $16\times 128$ for the second and third stages. This horizontally-elongated window configuration, while empirically motivated, demonstrates superior performance compared to conventional square attention windows, effectively capturing the inherent horizontal structure of range-projected LiDAR data.
Decoder: While the outputs of the encoder are multiscale features $F_{1},F_{2},F_{3},F_{4}$ with spatial size,
$\left[(H,W),\left(\frac{H}{2},\frac{W}{2}\right),\left(\frac{H}{4},\frac{W}{4}\right),\left(\frac{H}{8},\frac{W}{8}\right)\right]$ (8)
Table 1: Comparisons among different Hiera backbones on the test set of SemanticKITTI [5]. Interestingly, the SAM2-tiny model outperforms its larger counterparts despite having fewer parameters, demonstrating that model capacity does not necessarily correlate with performance in this task domain.
Backbone Params(M) mIoU car bicy moto truc o.veh ped b.list m.list road park walk o.gro build fenc veg trun terr pole sign
[…]
Table 2: Experiments with and without training augmentations. We observe a strong performance gain when introducing the proposed training augmentations from [41]. Comparisons among SAM2-tiny on SemanticKITTI test set [5].
Augs. mIoU car bicy moto truc o.veh ped b.list m.list road park walk o.gro build fenc veg trun terr pole sign
[…]
Table 3: Experiments with and without transfer-learning techniques using Cityscapes dataset [20]. We observe that overall our model does not benefit from extensive transfer-learning strategies introduced by [3]. Comparisons among SAM2-tiny on SemanticKITTI [5] test set.
Pretrain mIoU car bicy moto truc o.veh ped b.list m.list road park walk o.gro build fenc veg trun terr pole sign
[…]
Table 4: Comparisons among state-of-the-art LiDAR range view semantic segmentation approaches on the validation sequence of SemanticKITTI [5]. We compare our best performing model using the SAM2 with the tiny Hiera backbone.
Method VFM mIoU car bicy moto truc o.veh ped b.list m.list road park walk o.gro build fenc veg trun terr pole sign
RangeNet++[55] (IROS19) $\times$ 52.2 91.4 25.7 34.4 25.7 23.0 38.3 38.8 4.8 91.8 65.0 75.2 27.8 87.4 58.6 80.5 55.1 64.6 47.9 55.9
SqSegV3[99] (ECCV20) $\times$ 55.9 92.5 38.7 36.5 29.6 33.0 45.6 46.2 20.1 91.7 63.4 74.8 26.4 89.0 59.4 82.0 58.7 65.4 49.6 58.9
3D-MiniNet[2] (IROS20) $\times$ 55.8 90.5 42.3 42.1 28.5 29.4 47.8 44.1 14.5 91.6 64.2 74.5 25.4 89.4 60.8 82.8 60.8 66.7 48.0 56.6
SalsaNext[21] (ISVC20) $\times$ 59.5 91.9 48.3 38.6 38.9 31.9 60.2 59.0 19.4 91.7 63.7 75.8 29.1 90.2 64.2 81.8 63.6 66.5 54.3 62.1
[…]
MPF[1] (WACV21) $\times$ 55.5 93.4 30.2 38.3 26.1 28.5 48.1 46.1 18.1 90.6 62.3 74.5 30.6 88.5 59.7 83.5 59.7 69.2 49.7 58.1
FIDNet[112] (IROS21) $\times$ 59.5 93.9 54.7 48.9 27.6 23.9 62.3 59.8 23.7 90.6 59.1 75.8 26.7 88.9 60.5 84.5 64.4 69.0 53.3 62.8
[…]
RangeSAM(Ours) 60.9 91.7 42.0 44.3 41.2 36.1 55.9 54.4 30.2 92.2 67.5 77.2 28.6 89.8 62.8 84.5 66.9 69.8 58.9 62.4
and channel dimensions $[96,192,384,768]$. Following [98], we standardize the output channel dimensions to 256 using a modified Receptive Field Block [46] architecture that substitutes LayerNorm and GELU activations for the conventional BatchNorm [37] and ReLU, thereby achieving better compatibility with contemporary transformer architectures. Consistent with [27], we concatenate these four normalized feature maps and progressively reduce the dimensionality to match the number of target classes, while incorporating auxiliary classification heads at corresponding feature levels to enhance gradient flow during training.
Postprocessing: Dense evaluation on datasets like SemanticKITTI requires label propagation from processed points to the full-resolution point cloud via kNNk\-NN interpolation with majority voting. We use $k=7$, balancing efficiency and noise robustness within the typical range of 3-7 [15, 87, 3, 41].
[…]
where $\lambda_{i}$ represent empirically determined weighting coefficients. In our experimental configuration, we set $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}=1$, assigning equal importance to each loss component.
4 Dataset and implementation details
We now describe the datasets used, the training configurations applied, and the evaluation metrics employed.
4.1 Dataset selection and metrics:
We select the SemanticKITTI [5] and nuScenes [9] datasets as our primary training datasets. NuScenes [9] is a large-scale autonomous driving dataset containing 1,000 scenes of 20-second duration collected in urban settings across Boston and Singapore. The dataset captures diverse environmental conditions including varying illumination and weather patterns. Each scene is recorded using a 32-beam LiDAR system and annotated with 16 semantic categories. The standard split provides 28,130 point cloud frames for training and 6,019 frames for validation.
SemanticKITTI [5] is derived from the KITTI Vision Odometry Benchmark [27] and features urban driving scenarios captured in Germany. The dataset follows a sequence-based split: sequences 00-07 and 09-10 serve as the training set (19,130 scans), while sequence 08 is designated for validation (4,071 scans). The test set comprises sequences 11-21, containing 20,351 scans. Point cloud data is acquired using a 64-beam LiDAR sensor, with semantic annotations provided for 19 distinct classes. To maintain consistency across training and validation procedures, we standardize the nuScenes dataset by converting it to the SemanticKITTI format. We finally evaluate our experiments on the validation split of the SemanticKITTI benchmark dataset, we report performance using mean Intersection over Union (mIoU) [24] as the primary metric.
4.2 Data Augmentations:
Comprising global rotation, coordinate perturbation (jittering), geometric transformations (flipping), and random point elimination were employed to improve model generalization capabilities and enhance robustness to input variations we employ these with a probability of 1.0. Further we utilize the [41] introduced range-view-specific augmentation techniques that generate pseudo-images for direct application of mixing, union, shifting, and copy-paste operations with a probability to happen $[0.9,0.1,0.9,1.0]$.
4.3 Training strategy:
All experiments are done on 8 40GB A100 GPUs. We employed a per-GPU batch size of 2, yielding an effective global batch size of 16 across the distributed setup. To stabilize training, a linear warm-up of 5 epochs is applied during the first epoch, followed by a cosine annealing schedule [28] for gradual learning-rate decay. Recent studies [41, 3, 95, 96] have demonstrated that transformer-based architectures for point cloud processing exhibit significant performance gains when leveraging pretraining and transfer learning methodologies. Building upon the training protocols established in [41, 3], we adopt an extensive pretraining strategy on the nuScenes dataset [9]. Our proposed training framework consists of the following components: To ensure dataset compatibility and enable unified training, we converted the nuScenes dataset to the SemanticKITTI format. For our experiments, we employed the SAM2-tiny backbone as the base architecture and trained the model for 60 epochs on the unified dataset. Empirical evaluation revealed that utilizing heavier SAM2 backbone variants did not yield statistically significant performance improvements (table 1), while substantially increasing computational overhead and per-frame inference latency due to the increased parameter count. This finding guided our selection of the lightweight SAM2-tiny variant to optimize the trade-off between model performance and computational efficiency. The final model architecture comprises approximately 63 million parameters, providing an optimal balance between representational capacity and computational requirements for real-time applications.
Optimization: We use the AdamW optimizer with two different schemes: 1) Hiera backbone is trained with learning rate $lr=0.0004$ and the weight decay $\alpha=0.001$, 2) The rest of the model is trained with learning rate $lr=0.001$ and the weight decay $\alpha=0.0001$.
4.4 Evaluation
We present the experimental results obtained using our proposed RangeSAM methodology. Table 4 presents quantitative results obtained on the validation sequence 08 of the SemanticKITTI dataset [5]. Tables 1, 2, and 3 report performance metrics evaluated on the official SemanticKITTI test set [5]. Our approach demonstrates competitive performance with the SemanticKITTI benchmark (see Table 4), achieving results that are comparable to state-of-the-art methods in the field.
Figure 5: Qualitative segmentation examples of increasing difficulty using our model. (a) Urban intersection, (b) Suburban environment and (c) Highly cluttered scene.
4.5 Quantitative evaluation
Table 1 presents a comparative analysis of different SAM2 Hiera backbones. Due to computational constraints, our evaluation is limited to SAM2 small and tiny variants. The results demonstrate comparable mIoU performance between the Hiera tiny and small backbones, with the tiny variant offering computational advantages through reduced parameter count, resulting in faster inference times. Table 4 compares our model’s IoU performance against established baseline methods. While recent state-of-the-art architectures achieve mean IoUs of $65–75\%$ our model reaches $60.9\%$. This result shows that we successfully capture the most prevalent semantic classes, even though we underperform on rarer or smaller objects. Moreover, unlike other methods that do not use any vision foundation models, our approach integrates the SAM2 model [64], the current state of the art in semantic segmentation, demonstrating its effectiveness despite the modest mIoU gain. The quantitative results presented in Table 4 can be summarized as follows:
  • Several classes achieve IoUs above $80–90\%$. These are cars, road, building, and vegetation, which are both frequent in the dataset and relatively large structures in the LiDAR scans. The model does not lag far behind the state-of-the-art approaches.
  • Many classes fall in the $60–70\%$ IoU range. These are typically mid-frequency categories, such as trucks, fences, or terrain. Despite being the rarer classes, the model still managed to capture the details and once again, did not lag far behind the best approaches.
  • Several classes show very low IoUs in the range $29–47\%$. These correspond to rare and small objects in SemanticKITTI, such as motorcycles, bicycles, persons, trucks or other-vehicles. These are the rare and small object classes that are hard to capture, even with the modern approaches. This is where the model had alot of problems learning and interpolating.
Although our model delivers strong accuracy on common and semi-rare classes, its lower performance on long-tail categories reflects the limited training examples for these less frequent classes relative to our 63 M-parameter architecture. Nevertheless, our results remain competitive with leading baselines and show solid generalization to all well-represented categories. We are confident that expanding the dataset and further refining the design will help close the gap on tail classes and drive the model toward new state-of-the-art results.
4.6 Qualitative evaluation
Figure 5 illustrates three segmentation scenarios based on difficulty to segment it:
  • (a) Urban intersection: The model accurately delineates vehicles at a city crossroads as well as static elements such as trees, walls, fences and surrounding cars. Fine structures (e.g., fence posts) are also well localized.
  • (b) Suburban environment: Dominant classes like cars, vegetation and fences are correctly segmented. The model struggles with transmission towers, which were not represented in the training set, highlighting the challenge of unseen classes.
  • (c) Highly cluttered scene: Despite many overlapping objects (trees, fences, poles, signs), the model captures the major classes. Predictions for small or distant structures (signs, thin branches, far-off ground) exhibit noise and instability, but overall segmentation remains robust given the scene’s complexity.
4.7 Ablation Study
We conduct a comprehensive ablation study to systematically evaluate the impact of different training strategies on model performance.
Transfer learning: We followed a transfer learning approach established by other baseline methods[3] in this area of research. For the experiments in Table 3, we pretrained our model on the Cityscapes [20] dataset for 25 epochs to enhance feature learning before training on the target dataset. However, this approach led to reduced performance. We believe this may be due to the SAM2 model being extensively pretrained on large-scale image datasets, which could have caused a domain mismatch affecting transfer learning. Further investigation is needed to confirm this hypothesis and explore other adaptation strategies.
Range View Augmentations: The results (see Table 2) indicate that adding augmentations introduced by [41] to the training process improves performance across nearly all datasets, achieving a 10% gain in mIoU.
5 Conclusion
This work establishes the efficacy of VFMs for projection-base LiDAR point cloud semantic segmentation tasks. To the best of our knowledge, this is the first framework to leverage VFMs for range-view projected LiDAR point cloud segmentation. Despite the substantial domain gap between natural RGB images and range-view representations, our findings demonstrate that principled architectural adaptations tailored to the geometric properties of 3D point clouds can effectively bridge this modality shift and achieve competitive performance. Our contribution is a novel adaptation of the SAM2 architecture optimized for range-view point cloud data through targeted modifications in encoder and decoder. We develop a comprehensive training methodology that systematically integrates augmentation strategies and training protocols from prior works [15, 41, 95, 3]. Our empirical analysis reveals that extensive pretraining on 2D datasets, as proposed in [3], is unnecessary when employing visual foundation models (VFMs). Instead, we demonstrate that multi-dataset training across diverse 3D benchmarks yields superior performance gains. To facilitate reproducibility and future research, we will release both source code and model weights.
Future Work: A primary limitation of our current approach is computational complexity, particularly regarding parameter count. The convolution-free hiera backbone [70] presents inherent challenges for point cloud segmentation due to data sparsity, which necessitates convolutional operations to introduce essential locality bias. While our experiments demonstrate that RFBs[46] are critical for achieving competitive performance, they constitute the primary computational bottleneck. This parameter overhead currently precludes deployment in real-time applications, representing a key area for future optimization efforts.